1275:25::1411:x       give me right answers sir ji ...it is a reasoning queastion …....

Grade:12th pass

1 Answers

Aditya Gupta
4 years ago
by definition of ratios and proportion,
1275/25= 1411/x
or x= 1411*25/1275= 1411/51= 83/3.
so logically ans should be 83/3 or 26.667 but since it is a reasoning ques, ans may b different, and hence the question can be correctly solved if and only if options are provided. so kindly repost the ques along with all the options given.
kindly approve :)